Solve the equation log_3 (log_2 x) = 2.
Answer by math_helper(2461)


log_3(log_2(x)) = 2
log_3(n) = 2 ==> n=3^2=9
log_2(n) = 9 ==> n=2^9=512

Ans: x = 512

Chk: log_3(log_2(512)) = log_3(log_2(2^9)) = log_3(9) = log_3(3^2) = 2.